OpenStudy (anonymous):

simplify 110% into a fraction.

OpenStudy (anonymous):

110/1000 or 11/100.

OpenStudy (tkhunny):

100% = 100/100 50% = 50/100 900% = 900/100 2% = 2/100 What say you of 110%?